OpenStudy (anonymous):

y-5 over 3 = 1 -2 8 18 6

OpenStudy (mathstudent55):



OpenStudy (anonymous):

Um-so now you have plain y-5 =3?

OpenStudy (mathstudent55):

Correct. y - 5 = 3 Now you want y by itself, so what do you do?

OpenStudy (anonymous):

You subtract both sides by 5.

OpenStudy (anonymous):

-2 so answer choice a

OpenStudy (mathstudent55):

Wait. 5 is already being subtracted from y. To get rid off the 5, you need to add 5 to the left side becasue y - 5 + 5 = y Since you need to add 5 to the left side, you must also add 5 to the right side.

OpenStudy (anonymous):

ooooooooooooooh--NOW do you see what I mean? I RUSH things! Gragh. Ok-it would be y + 10 + 10?

OpenStudy (mathstudent55):

Not like that. Get back to theis step that you had correct: y - 5 = 3 Since 5 is being subtracted from y and you want y by itself, you need to add 5 to the left side. But whatever you do to the one side of the equation, you need to do the same to the other side, so add 5 to both sides. y - 5 + 5 = 3 + 5 y = 8

OpenStudy (anonymous):

Oh. Now I feel like a dummy. Thanks! I understand it now. Next question. 6x + 29 = 5 -4 -18 204 -144

OpenStudy (mathstudent55):

Follow each step I tell you. First, ask yourself, if the x on the left side were a number and you had to do the operations on the left sides, in what order would you do it in? Let's say instead of x you have 5. 6(5) + 29 How do you start that? What operation do you do first?

OpenStudy (anonymous):

Oh sorry I was walking my dog--dads orders. lol. OK-you would multiply the 6 and 5. then you would add the 29 to it.

OpenStudy (mathstudent55):

You need to undo what is being done to x, by doing the opposite opperations in the opposite order. 6x + 29 means multiply x by 6 first, then add that result to 29 Since adding 29 is done last, you must start by undoing that first using the opposite operation. the opposite of addition is subtraction. So you first must subtract 29, but whatever you do to one side of an equation you must do to the other side also. So first steop is subtract 29 from both sides. 6x + 29 = 5 6x + 29 - 29 = 5 - 29 6x = -24 Now the first operation before was to multiply 6 by x, so you must undo that last by dividing by 6. But you must do it to both sides, so 6x/6 = -24/6 x = -4 That's the final answer

OpenStudy (anonymous):

Oh. Wow. Thats an easier way than what my lesson told me to do. thanks!! next Q: 9a – b – 2a – 10b

OpenStudy (anonymous):

You have to simplify the expression answer choices: –7a + 11b 11a + 9b –11a – 9b 7a – 11b

OpenStudy (mathstudent55):

Combine terms with a together, and combine terms with b together. 9a - b - 2a - 10b = 9a - 2a - b - 10b Can you continue?

OpenStudy (anonymous):

I think I have a FAINT grasp on it-so you should probably go ahead. and yes of course you can go after this!! whenever u have to go---just tell me and go--i dont wanna hold you up!! thanks SO Much again!!!!!!!

OpenStudy (mathstudent55):

Remember above with this problem, combine a's together and combine b's together. a's and b's don't mix. They're like apples and oranges.

OpenStudy (mathstudent55):

9a - 2a - b - 10b = 7a - 11b Final answer

OpenStudy (anonymous):

ok thanks so much!

OpenStudy (anonymous):

I will close this Q and open another one up--and you can go now! :) Thanks again for your help! :D